Insert parentheses () to make each sentence true, then work it out underneath to Show your work. NOTE: you may need more than one set of parentheses 13, 1+6x8-2-5=23 14. 4 x 8 12 1 3 29 15. 6 x 4+8-2+4 74 16, 8-8x8+8+8=8
Answer #1

Answer :

13 ) To get the result of 23 insert the parenthesis as shown below

(((1 + 6 ) x 8) ÷ 2 ) - 5  

= ((( 7 x 8 )) ÷ 2 ) - 5

= ((56) ÷ 2) - 5

= 28 - 5

=23

14)   To get the result of 29 insert the parenthesis as shown below

(4 x 8) - (12 ÷(1 + 3))

= 32 - (12 ÷ 4)

= 32 - 3

= 29

Therefore, (4 x 8) - (12 ÷(1 + 3)) = 29

15) To get the result of 74 insert the parenthesis as shown below

(6 x( 4 + 8)) - 2 + 4

= (6 x 12 ) + 2

= 72 + 2

74

Therefore, (6 x( 4 + 8)) - 2 + 4 = 74

16) To get the result of 8 insert the parenthesis as shown below

(((8 - 8) x 8) ÷ 8 )+ 8

= ((0 x 8 ) ÷ 8) + 8

= (0 ÷ 8) + 8

= 0 + 8

=8

Therefore, (((8 - 8) x 8) ÷ 8 )+ 8 = 8